OpenStudy (anonymous):

Franklin left a $5 tip for his waitress after treating his friend to dinner. If the food bill was 31.25 , what was Franklin's tip as a percent?

OpenStudy (anonymous):

divide the amount (5) by the total (31.25)

OpenStudy (anonymous):

\[\frac{5}{31.25}\]

OpenStudy (anonymous):

make a equation TIP=x(31.25) where x is the percentage of tip 5=x(31.25) divide by 31.25 and you have your answer

OpenStudy (anonymous):

that will give you in the form of decimal but then you have to move the decimal two spots to the right and place a percentage sign

OpenStudy (anonymous):

multiply by 100 to get your decimal into percent form

OpenStudy (anonymous):

i got 0.16

OpenStudy (anonymous):

multiply by 100

OpenStudy (anonymous):

then place a percentage sign in the end

OpenStudy (anonymous):

thanks romero :)

OpenStudy (anonymous):

yw